# MS-0000-0000.00

### v1

@@ 3.1 Human Perception SA of Intelligent Agent
Early studies on the topic of human-AI teaming have documented a few benefits. AI or machines, in general, possess computational powers and might yield information and alternatives that humans do not consider on their own (Meyer et al. 2021). In particular, when AI serves the role of assistants, as humans are flexible and adaptable to changing contexts, humans can learn from the machines that can possibly play an instructional role in improving the quality or accuracy of humans’ decision-making (Choi et al. 2022), and overall welfare gains (Kleinberg et al. 2018). Humans might also welcome ceding control of certain decisions to agents. For example, humans may view algorithm-based task assignments as fairer than human-initiated task assignments (Bai et al. 2022). The ability of agents to process massive amounts of data for optimization is a contributor to the enhanced perception of SA.
Effective integration of human intelligence into agentic processes is still in its infancy. Humans possess distinct strengths that can contribute to better AI perception SA. The first strength is humans’ domain expertise. In decision-making, individuals, especially those with years of experience, can draw on their institutional knowledge from multiple, often undocumented, sources – commonly referred to as tacit knowledge (Cao et al. 2021, Choudhury et al. 2020, Kumar and Sharma 2022). Intrinsic expertise allows humans to potentially outperform AI or algorithms on perception SA. Second, humans possess sophisticated emotional intelligence, knowledge of institutional politics, and industry insights, which are still difficult for AI to master (Ibrahim et al. 2021). While humans can flexibly draw on these ‘private’ sources of information to inform their decisions, this type of data is not available or feasible to use in training AI models. Third, another key strength of humans is their interpersonal communication skills (Brynjolfsson and Mitchell 2017, Luo et al. 2021).
AI technologies present comparative advantages in routine-based, codifiable, or scripted tasks (Autor 2015). There are, however, contexts where AI can contribute to better emotional responses as well. High-volume tasks create cognitive overload on humans that places inherent limits on perception SA. Human ability to leverage intrinsic knowledge and emotional intelligence can be severely constrained during stressful situations, especially when multiple information sources have to be synthesized rapidly for decision-making. In such situations, AI can be more effective in formulating responses and augment human’s interpersonal communication skills (Luo et al. 2021). Effective human-AI teaming can complement each other’s strengths to create an optimized perception SA. For example, human-AI teams can effectively combine tacit knowledge, intuition, and empathy with a rapid synthesis of codifiable and unstructured knowledge and information (Feuerriegel et al. 2022, Lou and Wu 2021).
@@ 3.2 Human Comprehension SA of Intelligent Agent
Second, we ideally expect humans to correct AI’s mistakes in a timely manner, especially when facing ambiguous decision contexts. Studies show that when uncertain about their own decisions, especially when these conflict with AI’s recommendations, humans actively gather more information, search for alternative cues, or employ heuristic reasoning (Jussupow et al. 2021). However, this rethinking process is not guaranteed and depends on certain contextual factors, such as the presence of interesting and complex problems. Strategic rethinking requires a significant investment in cognitive resources, particularly when humans feel jaded performing routine work (Lebovitz et al. 2022). Given the interchangeable roles of AI and humans in completing similar tasks, humans may find it easier to apply cognitive abilities beyond their typical domain expertise. Effective utilization of human-AI collaboration can thus generate new, sustainable sources of competitive advantage (Krakowski et al. 2022).
@@ 3.3 Tasks in Human-AI Collaboration
As discussed in Section 2.3, group tasks typically encompass four key types, yet existing literature predominantly focuses on the decision-making phase, where humans and AI collaborate to solve specific problems. This narrow focus overlooks other types of group tasks, such as negotiation, communication, and coordination tasks. In traditional human-AI collaboration, AI often takes on static, predefined roles as an assistant, primarily executing human assigned tasks. This dynamic reflects an imbalance, as humans retain decision-making authority while AI lacks agency, leading to an unequal partnership. On the one hand, when humans and AI choose solutions for the same tasks, the literature suggests and confirms the importance of keeping humans in the decision-making loop to enhance effectiveness. Human workers are more likely to actively accept AI assistance or even make contributions if they could manipulate or make adjustments to AI recommendations (Dietvorst et al. 2018, Kawaguchi 2021). Additionally, researchers have examined how to incorporate human opinions (e.g., with voting strategy (Kamar et al. 2012), through a reinforcement learning framework (Koster et al. 2022)), when and how much to include human interventions (De et al. 2020, Lage et al. 2018), and even account for potential side effects of involving humans (e.g., generalizing rejection learning to mitigate biases held by external decision makers (Madras et al. 2018)).
@@
The emergence of generative AI systems challenges traditional work allocations in human-AI collaboration, often emphasizing the complementary roles of humans and AI. Earlier studies predominantly framed AI as excelling in cognitive intelligence tasks while viewing humans as better suited for creative, intuitive, or strategic roles (Jia et al. 2023). Generative AI, however, blurs these distinctions by demonstrating substantial creative capabilities (Toner-Rodgers 2024, Chen and Chan 2024). Human-AI chaining of tasks will, therefore, need to be reassessed to account for generative AI’s ability to contribute ideas, adapt to context, and collaborate more equitably.
@@ 3.4 Human-human Teaming
Human characteristics significantly influence shared SA and the ability to project teammates’ actions in human-human teaming. Extant research highlights that individual traits—such as trust, empathy, and communication skills—are crucial for cultivating and sustaining team SA. For instance, Salas et al. (2017) emphasize that trust among team members enhances information sharing and reduces hesitation in communication, leading to a more cohesive and synchronized awareness across the team. Additionally, Woolley et al. (2010) find that collective intelligence—linked to social sensitivity, balanced communication, and gender composition—strongly predicts group performance. Empathy and anticipation also play critical roles in the ability to project future actions of teammates. Studies suggest that when team members have a strong understanding of each other’s skills and behavioral tendencies, they can make more accurate predictions about how teammates will react under different circumstances (Entin and Serfaty 1999). This ability to project is especially crucial in high-stakes environments, where split-second decisions rely on accurately anticipating teammates’ responses. However, the literature also points to potential challenges; for instance, personality traits such as overconfidence or reluctance to share information can hinder team SA and the accuracy of projections. Overconfident individuals may underestimate the need for feedback, leading to gaps in SA, while those less inclined to communicate openly may limit the ability of the team to form accurate mental models (Cooke et al. 2017).
Coordination mechanisms in human-human teaming are essential for optimizing collaboration and ensuring efficient task execution. A well-structured workflow outlines task organization, responsibility distribution, and interaction sequences. Information, decision, and incentive structures are additional coordination mechanisms that can enhance workflow performance. With enhanced information visibility, adaptive workflows can be designed to create flexibility in task execution. Decision structures that clarify team member responsibilities are increasingly recognized as vital to operating in dynamic environments (Entin and Serfaty 1999).
Teams handling complex tasks require multiple agents with distinct skills, perspectives, and responsibilities. Designed for creative problem-solving, such teams face challenges in coordination and communication. Rainey and Tolk (2015) emphasize the importance of understanding team dynamics in multi-agent systems, suggesting that complexity arises not only from the individual characteristics of team members but also from their interactions and interdependencies. Teams performing complex tasks benefit from diversity in team member profiles (skills, academic background, experiences, socio-economic status, and other factors). For example,
@@ 4 Outlook for Future Research in a New Era of AI Renaissance
As AI-based intelligent agents continue to advance into increasingly proficient collaborators (USAF 2015), human-AI teaming exemplifies a paradigm where humans and intelligent agents engage in iterative, bidirectional, and multi-modal communications to pursue shared goals. Key phases of this collaboration include planning, facilitation, iteration, and testing. Maintaining optimal teaming dynamics in such settings demands not only coordination strategies but also conflict management to mitigate potential frictions and ensure effective task execution.
@@
We discuss the four research streams by first outlining the desired outcomes for effective human-AI teaming as a guiding objective. We then explore the associated positive and negative aspects (i.e., tensions), addressing key issues, measurement challenges, and mechanisms and solutions needed to resolve potential obstacles in human-AI teaming. In general, we adopt a cautiously optimistic approach to motivating human-AI teaming, avoiding overly positive or negative expectations on AI capabilities (Moser et al. 2024). Figure 2 provides a visual summary of the proposed research directions inspired by the extended team SA model, highlighting the significance of AI’s contextual understanding and the team’s adaptive capabilities.
@@ 4.1 Understanding Human Learning: Human Projection of AI Agents
Projection in team SA involves anticipating the future actions and consequences of both human and intelligent agent team members. As AI-based intelligent agents evolve, their increasing autonomy and adaptability necessitate a deeper understanding of how humans adjust their expectations and behaviors to these advanced collaborators.
@@ 4.1.1   Expected Outcomes
AI is expected to become an active contributor/member rather than a passive tool in team (Vanneste and Puranam 2024, Dennis et al. 2023). This shift envisions AI agents to process complex information and develop actionable insights to complement human creativity and decision-making. AI and human partnership can enhance problem-solving efficiency, reduce the cognitive workload on human team members, and boost productivity. Particularly, Salas et al. (2017) emphasize that trust among team members is essential to enhancing information sharing and reducing communication hesitancy. If trust in the AI team collaborator increases, it would lead to a more cohesive team dynamic. Trust in AI is expected to grow as its contributions become more accurate and contextually relevant, enabling smoother workflows and more confident task delegation. However, this integration brings challenges. Over-reliance on AI may suppress human creativity, learning, and critical thinking, as individuals increasingly defer to AI-generated solutions without adequate investigation (Fügener et al. 2021, Dell’Acqua et al. 2023). Misaligned expectations, where AI outputs fail to meet team goals, can lead to frustration, diminished trust, and resistance to adoption. Furthermore, excessive dependence on AI for routine and complex tasks risks eroding human expertise, leaving teams vulnerable in scenarios where AI is unavailable or unreliable. Addressing these tensions is crucial for fostering sustainable and balanced collaboration in human-AI teams.
@@ 4.1.2   Evaluations and Issues
Altering Human Expectations:  As discussed in Section 2.1, intelligent agents today are capable of interactions that are much strongly anchored on the task context. Interactions over time at varying levels of intensity in task contexts will contribute to expectation formation for the human collaborator. Expectations represent a form of human learning that relates to human projection (Level 3) toward the external situation of the intelligent agent and the dynamics of teamwork (Frydman 1982). In iterative interactions, humans may change their expectations about the next input from their AI teammate, which is much more nuanced and complex than what has been studied in the literature on technology usage. The rich conversational interaction between AI and humans can lead to anticipative reasoning, planning, and discovery in both human and AI team members. Such a symbiotic interaction can result in better teamwork (Bayer and Renou 2024). The new interaction styles can radically alter human evaluations of technology use where AI collaborations are not merely the cumulative or average result of individual stages of the interaction process as it is usually studied, but can often disproportionately be influenced by the most salient (“peak”) and final (“end”) moments of interactions (i.e., peak-end rule) (Fredrickson and Kahneman 1993). For example, an AI-generated output during a critical task phase or a resolution provided at the conclusion of an interaction can significantly enhance human expectations and satisfaction, even if earlier stages were less effective. Studying such complex interaction processes and outcomes poses an opportunity for IS researchers. Methodologically, constructs for measuring the dynamics of interactions are either nonexistent in AI research or underused. It is unclear if measurements used in human-human teaming contexts would fit well in examining human-AI teaming. Further, a qualitative understanding of the learning process at the dyadic level presents new grounded theory development opportunities. Future research can also take mixed methods approaches to describe the dynamic formation of expectations toward (1) AI outputs and reasoning, (2) cognitive loads, and (3) task and team performance. Learning within the human-AI dyad can also have an impact on human-human interactions within the team. Differing trust levels of AI among team members, task-AI fit across different team responsibilities, and the associated feedback loops present a complex challenging research domain that is an opportunity for new theory development in IS. Accumulated theoretical knowledge in this domain can help organizations redesign tasks, calibrate task assignments, and reconfigure workflows.
@@
Quantification of Human Responses:  Following the canonical expectation-confirmation framework that is used to understand information systems continuance (Bhattacherjee 2001), this kind of focal specific human responses relates to human projection (Level 3) toward external situation. The augmented AI capability may reshape humans’ attitudes and learning behaviors (Lu et al. 2024). It is imperative to further disambiguate human-AI interactions at the task level to gain new insights on how humans lean on the extended cognitive support for their work. Human responses can differ within a task as it pertains to idea generation, solution selection, conflict resolution, and task execution. In particular, incorporating the peak-end rule discussed earlier, it becomes critical to investigate how humans react to updates and augmentations provided by AI during task execution, particularly at moments of high significance or at the conclusion of interactions under each task type. These moments could have an outsized impact on humans’ response to AI. Key metrics include response times, decision accuracy, and reliance on AI suggestions. Longitudinal studies can track changes in human (1) attitudes and (2) adaptation behavior toward AI over time, offering insights into the evolving trust and collaboration dynamics between humans and AI.

### v2

@@ 3.1 Human Perception SA of Intelligent Agent
Early studies on the topic of human-AI teaming have documented a few benefits. AI or machines, in general, possess computational powers and might yield information and alternatives that humans do not consider on their own (Meyer et al. 2021). In particular, when AI serves the role of assistants, as humans are flexible and adaptable to changing contexts, humans can learn from the machines that can possibly play an instructional role in improving the quality or accuracy of humans’ decision-making (Choi et al. 2022), and overall welfare gains (Kleinberg et al. 2018). Humans might also welcome ceding control of certain decisions to AI. For example, humans may view algorithm-based task assignments as fairer than human-initiated task assignments (Bai et al. 2022). The ability of AI to process massive amounts of data for optimization is a contributor to the enhanced perception of SA.
Effective integration of human intelligence into agentic processes is still in its infancy. Humans possess distinct strengths that can contribute to better AI perception SA. The first strength is humans’ domain expertise. In decision-making, individuals, especially those with years of experience, can draw on their institutional knowledge from multiple, often undocumented, sources – commonly referred to as tacit knowledge (Cao et al. 2021, Choudhury et al. 2020, Kumar and Sharma 2022). Intrinsic expertise allows humans to potentially outperform AI on perception SA. Second, humans possess sophisticated emotional intelligence, knowledge of institutional politics, and industry insights, which are still difficult for AI to master (Ibrahim et al. 2021). While humans can flexibly draw on these ‘private’ sources of information to inform their decisions, this type of data is not available or feasible to use in training AI models. Third, another key strength of humans is their interpersonal communication skills (Brynjolfsson and Mitchell 2017, Luo et al. 2021).
AI presents comparative advantages in routine-based, codifiable, or scripted tasks (Autor 2015). There are, however, contexts where AI can contribute to better emotional responses as well. High-volume tasks create a cognitive overload on humans that places inherent limits on perception SA. Human ability to leverage intrinsic knowledge and emotional intelligence can be severely constrained during stressful situations, especially when multiple information sources have to be synthesized rapidly for decision-making. In such situations, AI can be more effective in formulating responses and augmenting human interpersonal communication skills (Luo et al. 2021). Effective human-AI teaming can complement each other’s strengths to create an optimized perception SA. For example, human-AI teaming can leverage tacit knowledge, intuition, and empathy with rapid synthesis of codifiable knowledge and unstructured information (Feuerriegel et al. 2022, Lou and Wu 2021).
@@ 3.2 Human Comprehension SA of Intelligent Agent
Second, we could ideally expect humans to correct AI’s mistakes in a timely manner, especially in ambiguous decision contexts. Prior studies show that when uncertain about their own decisions, especially when these conflict with AI’s recommendations, humans actively gather more information, search for alternative cues, or employ heuristic reasoning (Jussupow et al. 2021). However, this rethinking process is not guaranteed and depends on certain contextual factors, such as the presence of interesting and complex problems. Strategic rethinking requires a significant investment in cognitive resources, particularly when humans feel jaded performing routine work (Lebovitz et al. 2022). Given the interchangeable roles of AI and humans in completing similar tasks, humans may find it easier to apply cognitive abilities beyond their typical domain expertise. Effective utilization of human-AI collaboration can thus generate new, sustainable sources of competitive advantage (Krakowski et al. 2022).
@@ 3.3 Tasks in Human-AI Collaboration
As discussed in Section 2.3, group tasks typically encompass four key types, yet existing literature predominantly focuses on the decision-making phase, where humans and AI collaborate to solve specific problems. This narrow focus overlooks other types of group tasks, such as negotiation, communication, and coordination tasks. In traditional human-AI collaboration, AI often takes on static, predefined roles as an assistant, primarily executing human assigned tasks. This dynamic reflects an imbalance, as humans retain decision-making authority while AI lacks agency, leading to an unequal partnership. On the one hand, when humans and AI choose solutions for the same tasks, the literature suggests and confirms the importance of keeping humans in the decision-making loop to enhance effectiveness. Human workers are more likely to accept AI assistance or even make contributions if they can manipulate or make adjustments to AI recommendations (Dietvorst et al. 2018, Kawaguchi 2021). In addition, researchers have examined how to incorporate human opinions (e.g., with voting strategy (Kamar et al. 2012), through a reinforcement learning framework (Koster et al. 2022)), when and how much to include human interventions (De et al. 2020, Lage et al. 2018), and even account for potential side effects of involving humans (e.g., generalizing rejection learning to mitigate biases held by external decision makers (Madras et al. 2018)).
@@
The emergence of generative AI systems challenges traditional work allocations in human-AI collaboration, often emphasizing the complementary roles of humans and AI. Prior studies view humans as better suited for creative, intuitive, or strategic roles (Jia et al. 2023). Generative AI, however, blurs these distinctions by demonstrating substantial creative capabilities (Toner-Rodgers 2024, Chen and Chan 2024). Human-AI chaining of tasks will, therefore, need to be reassessed to account for generative AI’s ability to contribute ideas, adapt to context, and collaborate more equitably.
@@ 3.4 Human-human Teaming
Human characteristics play critical roles in shared SA and the ability to project teammates’ actions in human-human teaming. Extant research highlights that individual traits—such as trust, empathy, and communication skills—are crucial for cultivating and sustaining team SA. For example, Salas et al. (2017) demonstrate that trust among team members enhances information sharing and reduces hesitation in communication, leading to a more cohesive and synchronized awareness across the team. In addition, Woolley et al. (2010) find that collective intelligence—linked to social sensitivity, balanced communication, and gender composition—strongly predicts group performance. Empathy and anticipation further support the capacity to project future actions of teammates. Prior studies suggest that when team members have a strong understanding of each other’s skills and behavioral tendencies, they can make more accurate predictions about how teammates will react under different circumstances (Entin and Serfaty 1999). This ability to project is especially crucial in high-stakes environments, where split-second decisions rely on accurately anticipating teammates’ responses. However, the literature also points to potential challenges; for example, personality traits such as overconfidence or reluctance to share information can hinder team SA and the accuracy of projections. Overconfident individuals may underestimate the need for feedback, leading to gaps in SA, while those less inclined to communicate openly may limit the ability of the team to form accurate mental models (Cooke et al. 2017).
Coordination mechanisms in human-human teaming are also essential for optimizing collaboration and ensuring efficient task execution. A well-structured workflow outlines task organization, responsibility distribution, and interaction sequences. Information, decision, and incentive structures are additional coordination mechanisms that can enhance workflow performance. With enhanced information visibility, adaptive workflows can be designed to create flexibility in task execution. Decision structures that clarify team member responsibilities are increasingly recognized as vital to operating in dynamic environments (Entin and Serfaty 1999).
When teams are tasked with solving complex problems, coordination becomes more challenging due to the involvement of multiple agents with diverse skills, perspectives, and responsibilities. Such teams are often deliberately designed, and their success depends on navigating the increased complexity of communication and collaboration. Rainey and Tolk (2015) emphasize the importance of understanding team dynamics in multi-agent systems, suggesting that complexity arises not only from the individual characteristics of team members but also from their interactions and interdependencies. Teams performing complex tasks benefit from diversity in team member profiles (including skills, academic background, experiences, socio-economic status, and other factors).
@@ 4 Outlook for Future Research in a New Era of AI Renaissance
As AI agents continue to advance into increasingly proficient collaborators (USAF 2015), human-AI teaming exemplifies a paradigm where humans and agents engage in iterative and multi-modal communications to pursue shared goals. Key phases of this collaboration include planning, facilitation, iteration, and testing. Maintaining optimal teaming dynamics in such settings demands not only coordination strategies but also conflict management to mitigate potential frictions and ensure effective task execution.
@@
We structure the four research streams by first outlining the desired outcomes for effective human-AI teaming as a guiding objective. We then explore the associated positive and negative aspects (i.e., tensions), addressing key issues, measurement challenges, and mechanisms and solutions needed to resolve potential obstacles in human-AI teaming. In general, we adopt a cautiously optimistic approach to motivating human-AI teaming, avoiding overly positive or negative expectations on AI capabilities (Moser et al. 2024). Figure 2 provides a visual summary of the proposed research directions inspired by the extended team SA model, highlighting the significance of AI’s contextual understanding and the team’s adaptive capabilities.
@@ 4.1 Understanding Human Learning: Human Projection of AI Agents
Projection in team SA involves anticipating the future actions and consequences of both human and agent team members. As AI agents evolve, their increasing autonomy and adaptability necessitate a deeper understanding of how humans adjust their expectations and behaviors to these advanced collaborators.
@@ 4.1.1   Expected Outcomes
AI agents are expected to become active contributors/members rather than passive tools in team (Vanneste and Puranam 2024, Dennis et al. 2023). This shift envisions AI agents to process complex information and develop actionable insights to complement human creativity and decision-making. AI and human partnership can enhance problem-solving efficiency and reduce the cognitive workload on human team members. Particularly, Salas et al. (2017) emphasize that trust among team members is essential to enhancing information sharing and reducing communication hesitancy. If trust in the AI team collaborator increases, it would lead to a more cohesive team dynamic. Trust in AI is expected to grow as its contributions become more accurate and contextually relevant, enabling smoother workflows and more confident task delegation. However, this integration brings challenges. Over-reliance on AI may suppress human creativity, learning, and critical thinking, as individuals increasingly defer to AI-generated solutions without adequate investigation (Fügener et al. 2021, Dell’Acqua et al. 2023). Misaligned expectations, where AI outputs fail to meet team goals, can lead to frustration, diminished trust, and resistance to adoption. Furthermore, excessive dependence on AI for routine and complex tasks risks eroding human expertise, leaving teams vulnerable in scenarios where AI is unavailable or unreliable. Addressing these tensions is crucial for fostering sustainable and balanced collaboration in human-AI teams.
@@ 4.1.2   Evaluations and Issues
Altering Human Expectations:  As discussed in Section 2.1, intelligent agents today are capable of interactions that are much more strongly anchored on the task context. Interactions over time at varying levels of intensity in task contexts will contribute to expectation formation for the human collaborator. Expectations represent a form of human learning that relates to human projection (Level 3) toward the external situation of the intelligent agent and the dynamics of teamwork (Frydman 1982). In iterative interactions, humans may change their expectations about the next input from their AI teammate, which is much more nuanced and complex than what has been studied in the literature on technology usage. The rich conversational interaction between AI and humans can lead to anticipative reasoning, planning, and discovery in both human and AI team members. Such a symbiotic interaction can result in better teamwork (Bayer and Renou 2024). The new interaction styles can radically alter human evaluations of technology use where AI collaborations are not merely the cumulative or average result of individual stages of the interaction process as it is usually studied, but can often disproportionately be influenced by the most salient (“peak”) and final (“end”) moments of interactions (i.e., peak-end rule) (Fredrickson and Kahneman 1993). For example, an AI-generated output during a critical task phase or a resolution provided at the conclusion of an interaction can significantly enhance human expectations and satisfaction, even if earlier stages were less effective. Studying such complex interaction processes and outcomes poses an opportunity for IS researchers. Methodologically, constructs designed to capture the dynamics of interactions remain either insufficiently developed or underutilized in AI research. It is unclear if measurements used in human-human teaming contexts would fit well in examining human-AI teaming. Furthermore, a qualitative understanding of the learning process at the dyadic level presents new grounded theory development opportunities. Future research can take mixed-method approaches to describe the dynamic formation of expectations toward (1) AI outputs and reasoning, (2) cognitive loads, and (3) task and team performance. Learning within the human-AI dyad can also have an impact on human-human interactions within the team. Differing trust levels of AI among team members, task-AI fit across different team responsibilities, and the associated feedback loops present a complex challenging research domain that is an opportunity for new theory development in the information systems field. Accumulated theoretical knowledge in this domain can help organizations redesign tasks, calibrate task assignments, and reconfigure workflows.
@@
Quantification of Human Responses:  Following the canonical expectation-confirmation framework that is used to understand information systems continuance (Bhattacherjee 2001), this kind of focal specific human responses relates to human projection (Level 3) toward external situation. The augmented AI capability may reshape humans’ attitudes and learning behaviors (Lu et al. 2024). It is imperative to further disambiguate human-AI interactions at the task level to gain new insights into how humans lean on the extended cognitive support for their work. Human responses can differ within a task as it pertains to idea generation, solution selection, conflict resolution, and task execution. In particular, incorporating the peak-end rule discussed earlier, it becomes critical to investigate how humans react to updates and augmentations provided by AI during task execution, particularly at moments of high significance or at the conclusion of interactions under each task type. These moments could have an outsized impact on humans’ response to AI. Key metrics include response times, decision accuracy, and reliance on AI suggestions. Longitudinal studies can track changes in human (1) attitudes and (2) adaptation behavior toward AI over time, offering insights into the evolving trust and collaboration dynamics between humans and AI.